a tank is filled up to a height 2H WITH A LIQUID and is placed on a platform of height H from the ground .the distance x from the ground where a small hole is punched to get maximum range R is?      


 

Grade:12

6 Answers

abhinav anand
11 years ago

(squareroot of 24)H

Rishabh Maggirwar
11 years ago

The pressure of liquid is maximum at the bottom of the liquid in the tank.

So to get the maximum range, the hole must be pierced at the bottom of the tank, i.e., 

the required  height x must be equal to H.

 

with regards, 

Rishabh class xii

Abhishekh kumar sharma
11 years ago

the value of x= 3h/2

since v=√2gh   and in questiion h=3h-x

and range r=vt (projectile)

so    r= [√2g(3h-x)][√2x/g]

simplifying

we have r is maximum if (3h-x).x is maximum

which is maximum at x=3h/2

Akash Kumar Dutta
11 years ago

its the same question asked just before.

ans is x=3H/2.

gauhar singh
11 years ago

3H/2

DEV MUKHERJEE
11 years ago
let x be the distance fom the grond then here we have a horizontal projectlie time to come to grond =1/2 gt^2=x forwrd velocity =sqrt(2g(3h-x)) range=time *velocity; range =3h/2 (we would get a function for range which can be maximised using maxima and minima dr/dx=0 )